OpenStudy (anonymous):

if (a,b,c) is a geometric sequence and x is an arithmetic mean between a and b, and y is an arithmetic mean between b and c , prove that a/x +c/y=2

myininaya (myininaya):

Lets see we have \[\frac{a+b}{2}=x \text{ and } \frac{b+c}{2}=y\] We want to prove \[\frac{a}{x}+\frac{c}{y}=2\]

OpenStudy (anonymous):

yup

myininaya (myininaya):

We also have (a,b,c) is a geometric sequence..... That means a*m=b and b*m=c for some number m

myininaya (myininaya):

Thinking...

myininaya (myininaya):

What happens if we try to solve both those in my first post for b and set them equal what would we get...

myininaya (myininaya):

\[a+b=2x \text{ and } b+c=2y \] => \[b=2x-a \text{ and } b=2y-c\] => \[2x-a=2y-c\]

myininaya (myininaya):

Now can you solve this for 2?

myininaya (myininaya):

\[2x-2y=a-c\]

OpenStudy (anonymous):

still i tried that

myininaya (myininaya):

\[2(x-y)=a-c =>2 =\frac{a-c}{x-y}\] hmmm.....

OpenStudy (anonymous):

well u can say that a is a , b is ar and c is ar^2

OpenStudy (anonymous):

since its GS

OpenStudy (anonymous):

and plug in new values in both first equations then u get r= y/x

myininaya (myininaya):

True

OpenStudy (anonymous):

and since its GS so b is a geometric mean of a and c

OpenStudy (anonymous):

b^2=ac

myininaya (myininaya):

\[2=\frac{a-c}{x-y}\] \[2=\frac{a-ar^2}{x-xr} \text{ since } c=ar^2 \text{ & } y=xr\] So we have \[2=\frac{a(1-r^2)}{x(1-r)}\] \[2=\frac{a(1-r)(1+r)}{x(1-r)}\] \[2=\frac{a(1+r)}{x}\] \[2=\frac{a+ar}{x}\] \[2=\frac{a}{x}+\frac{ar}{x}\] guess what.........:)

myininaya (myininaya):

\[2=\frac{a}{x}+\frac{b}{x}\] oh i thought i had it

OpenStudy (anonymous):

haha but nice try keep going

myininaya (myininaya):

oh but wait

myininaya (myininaya):

we do have it

myininaya (myininaya):

\[\frac{c}{y}=\frac{ar^2}{xr}=\frac{ar}{x}\]

myininaya (myininaya):

right?

OpenStudy (anonymous):

yup

myininaya (myininaya):

\[2=\frac{a}{x}+\frac{ar}{x}=\frac{a}{x}+\frac{c}{y}\]

myininaya (myininaya):

Wow @hossam1122 I like this problem

myininaya (myininaya):

It is very cute.

OpenStudy (anonymous):

lol tnx :D

myininaya (myininaya):

do you have any questions? There were some things you mentioned that i didn't think of that we used to show this so really great job. I love it when askers want to engage. Thanks so much.

OpenStudy (anonymous):

you r welcome ;)